OpenStudy (anonymous):

Simplify 4 6/7 + 9 1/4 + 2 3/4 + 1/7.

hero (hero):

Well, to make things easier to do this: (4+1)+(6/7 + 1/7) +(9+2) + (1/4 + 3/4)

hero (hero):

And by doing that, it just so happens that (5) + (7/7) + (11) + (4/4) = 5 + 1 + 11 + 1 = 6 + 12 = 18